x-2y=6
3x-6y=18 how many solutions does this equation have

Answer by checkley77(12844)

x-2y=6 multiply by 3 & add.
3x-6y=18
3x-6y=18
------------
6x=36
x=36/6
x=6 ans.
6-2y=6
-2y=6-6
-2y=0
y=0 ans.
Proof:
3*6-6*0=18
18+0=18
18=18
One solution: (6,0).